hi i need help with these three problems
a) x^3e^x-4xe^x=0

b)e^1-5x= 3 square root of e
c) (lnx)^4= 4lnx

Answer by stanbon(48535)

a) x^3e^x-4xe^x=0
xe^x(x^2-4) = 0
xe^x = 0 or x^2-4 = 0
x = 0 or x=2 or x=-2
----------------------------
b)e^(1-5x) = 3 sqrt(e)
Take the natural log of both sides to get:
1-5x = ln3 + lne
5x = 1 - ln3 - 1
5x = -ln3
x = (1/5)(-1.0986..._
x = -0.219722...
Cheers,
Stan H.

c) (lnx)^4= 4lnx